Find the class boundaries, midpoints, and widths for each class.
a. 37 à ƒ ¢ € “ 49
b. 88 à ƒ ¢ € “ 105
c. 875 à ƒ ¢ € “ 899
d. 12.3 à ƒ ¢ € “ 15.5
e. 3.17 à ƒ ¢ € “ 5. 89
1
Expert's answer
2021-10-13T13:27:59-0400

a.37-49

Class boundaries = 37 - 0.5 and 49 + 0.5

So class boundaries = 36.5 and 49.5

Midpoint = (36.5 + 49.5)/2 = 43

Class width = 49.5 - 36.5 = 13


b.88-105

Class boundaries = 88 - 0.5 and 105 + 0.5

So class boundaries = 87.5 and 105.5

Midpoint = (87.5 + 105.5)/2 = 96.5

Class width = 105.5 - 87.5 = 18


c.875-899

Class boundaries = 875 - 0.5 and 899 + 0.5

So class boundaries = 874.5 and 899.5

Midpoint = (874.5 + 899.5)/2 = 887

Class width = 899.5 - 874.5 =25


d.12.3-15.5

Class boundaries = 12.3 - 0.05 and 15.5 + 0.05

So class boundaries = 12.25 and 15.55

Midpoint = (12.25 + 15.55)/2 = 13.9

Class width = 15.55 - 12.25 = 3.3


e.3.17-5.89

Class boundaries = 3.17-0.005 and 5.895 + 0.005

So class boundaries = 3.165 and 5.895

Midpoint = (3.165 + 5.895)/2 = 4.53

Class width = 5.895 - 3.165 = 2.73